Find two numbers whose sum is 92 if the first is four more than seven times the second.
Answer by edjones(7311)

x+y=92
x=7y+4
replace x with 7y+4 in the 1st equation. 7y+4+y=92
subtract 4 from each side. 8y=88
divide 8 into each side. y=11, x=7*11+4=81
Ed